40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3258     Entry time: Wed Jul 21 12:20:58 2010
Author: Razib 
Type: Update 
Category: Phase Camera 
Subject: Weekly update 

This past week I have worked on the following:

1. Setting up the infrastructure to do noise analysis: We added a temporary channel on the DAQ to connect to the PD 55 which we are using to take the power measurement. Before that, I connected the PD55 to an oscilloscope and recorded the power.

    phase_cam_setup_21_07_2010.jpg

The power at PD55 as measured using the oscilloscope = 600 µV.

Then I tried to calibrate the channel by sending up a signal from the function generator and measuring up the offset.. However, the channels seems noisy enough, especially due to electronics noise as suggested by the measurements and FFT calculation.

2. I worked on trying to sync the data acquisition of the PD and the CAM. After sometime spent on fiddling with the software method such as taking images at stamped time and then lining them up with the daq timestamps, I found a hardware method as suggested by Aidan. It was putting up a shutter (Uniblitz shutter and driver VMMD1) in the setup. I synced the shutter with the camera for which I had to tear apart the previously made trigger box and add a sync output from the camera (took a while as I also had to make a new cable).

3. I worked (still working) on making a differential amplifier to blow up the signal from the PD.

 

 

 

ELOG V3.1.3-